OpenStudy (anonymous):

solve (2x^3-4x^2+3)+(x^2-3x^2+1)

OpenStudy (anonymous):

add them! :)

OpenStudy (anonymous):

idk how to :C

OpenStudy (anonymous):

what is \(3+1\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

btw is it really \( (2x^3-4x^2+3)+(x^2-3x^2+1) \) or perhaps \( (2x^3-4x^2+3)+(x^3-3x^2+1) \)

OpenStudy (anonymous):

you see the difference? does the second one have two square terms \(x^2\) and \(-x^2\) or does it have a cube term \(x^3\) and a square term \(-2x^2\) ?

OpenStudy (anonymous):

is the first one that you posted

OpenStudy (anonymous):

i am going to guess it is \[(2x^3-4x^2+3)+(x^3-3x^2+1)\] am i right?

OpenStudy (anonymous):

yup thats it

OpenStudy (anonymous):

ok lets go slow (but not too slow)

OpenStudy (anonymous):

lol ok :)

OpenStudy (anonymous):

we already decided that \(3+1=4\) so that will be the number out at the end

OpenStudy (anonymous):

yup i got that part

OpenStudy (anonymous):

like terms have the same exponent, so we have like terms of \(x^3\) and \(2x^3\)

OpenStudy (anonymous):

and just like \(1000+2000=3000\) or one apple and two apples is three apples, we get \[x^3+2x^3=3x^3\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

so u dont add the exponents?

OpenStudy (anonymous):

hell no i was just about to say don't screw around with the exponents when you are adding that is for when you multiply

OpenStudy (anonymous):

ok so then what ?

OpenStudy (anonymous):

look at my first example \(1000+2000=3000\) i put it there on purpose, because \(1000=10^3\) and \(2000=2\times 10^3\) and \(10^3+2\times 10^3=3\times 10^3\) powers stay the same don't change them!!

OpenStudy (anonymous):

oh ok i got it

OpenStudy (anonymous):

so finally we get \(-4x^2-3x^2\)

OpenStudy (anonymous):

any guesses?

OpenStudy (anonymous):

-7x^2?

OpenStudy (anonymous):

you got it !

OpenStudy (anonymous):

final answer: \[\large 3x^3-7x^2+4\]

OpenStudy (anonymous):

omg thats what i got yay ok i got it i think i understand now C:

OpenStudy (anonymous):

that wasn't that bad, right?

OpenStudy (anonymous):

we can do another one if you like

OpenStudy (anonymous):

it seems easy with your help :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i got one

OpenStudy (anonymous):

\[(5a-4b+3c) + (6a-4b+2c)\] @satellite73

OpenStudy (anonymous):

ok like terms are the ones with the same letter

OpenStudy (anonymous):

you do this mostly with your eyes \[5a+6a=?\]

OpenStudy (anonymous):

11a

OpenStudy (anonymous):

k good, on to the next one \(-4b-4b=?\)

OpenStudy (anonymous):

-8b

OpenStudy (anonymous):

k next \(3c+2c=?\)

OpenStudy (anonymous):

omg okay so i got this one but idk how to do this other one you think we could move to the other one ? pleaseeee

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

okay i always have troublee with this (x+2) (x-2)

OpenStudy (anonymous):

you have to do four multiplications 1) \(x\times x=x^2\) 2) \(2\times x=2x\) 3) \(-2\times x=-2x\) 4) \(2\times -2=-4\)

OpenStudy (anonymous):

you get \[x^2+2x-2x-4=x^2-4\]

OpenStudy (anonymous):

oh omg now i get it aw thanks for all your help C:

OpenStudy (anonymous):

yw btw with practice you will see that \((a+b)(a-b)=a^2-b^2\) always, so \((x+2)(x-2)=x^2-4\) will come right away

OpenStudy (anonymous):

oh ok omg @satellite73 one more please please please it iwll be the last one i promise ?

OpenStudy (anonymous):

no problem, just ask

OpenStudy (anonymous):

ok so how would you solve this -2x^4(3x^2-5x)

OpenStudy (anonymous):

you mean multiply right?

OpenStudy (anonymous):

\[ -2x^4(3x^2-5x) \] \[=-2x^4\times 3x^2+2x^4\times 5x\] now this is where you add the exponents

OpenStudy (anonymous):

you should get \[-6x^6+10x^5\]

OpenStudy (anonymous):

-6x^6+10x^5

OpenStudy (anonymous):

yes you go it!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

not too bad...

OpenStudy (anonymous):

thanks c: i know im slow sorry :C